OpenStudy (anonymous):

Find an equation of the line containing the given pair of points. (2,1) and (6,4) Did I figure this out correctly?

OpenStudy (anonymous):

Find slope using the formula (y2-y1)/(x2-x1) Then use it along with one of the points to find eqn of line using slope point formula: y-y1=m(x-x1)

OpenStudy (anonymous):

I was posting the work so I could see if I did it correctly.

OpenStudy (anonymous):

sure go ahead

OpenStudy (anonymous):

m=\[4-1 \over 6-2\]

hero (hero):

I thought you were going to use y = mx + b

OpenStudy (anonymous):

Which would then be, 4-1 which is 3 and 6-2 which is 4 so the answer would be \[3 \over 4\]

OpenStudy (anonymous):

Lol, why you say that, Hero? I do need to be able to do it both ways right?

hero (hero):

Yes, of course :P

OpenStudy (anonymous):

yea it can be done both ways, if you want to use y=mx+b then you will have to find b by substituting one of the points for x and y

hero (hero):

No, @niravshah08, my way doesn't assume that you know m or b .

hero (hero):

I was saying to strictly use y = mx + b to solve the problem. No slope formula.

OpenStudy (anonymous):

ohh I misread it

OpenStudy (anonymous):

If I used y=mx+b after I found the slope out it would be like this: 4=3/4(6)+b which then would leave me with b = - 1/2

OpenStudy (anonymous):

it will be positive 1/2

OpenStudy (anonymous):

The answer would be the same if I plugged in the other numbers, which makes the lines parallel.

OpenStudy (anonymous):

Yeah, my finger hit the minus sign when I was typing lol. I type too fast sometimes.

OpenStudy (anonymous):

lol